### v1

# Model Editing at Scale leads to Gradual and Catastrophic Forgetting

Abstract
Editing knowledge in large language models is an attractive capability to have which allows us to correct incorrectly learnt facts during pre-training, as well as update the model with an ever-growing list of new facts. While existing model editing techniques have shown promise, they are usually evaluated using metrics for reliability, specificity and generalization over one or few edits. We argue that for model editing to have practical utility, we must be able to make multiple edits to the same model. With this in mind, we evaluate the current model editing methods at scale, focusing on two state of the art methods - ROME and MEMIT. We find that as the model is edited sequentially with multiple facts, it continually forgets previously edited facts and the ability to perform downstream tasks. This forgetting happens in two phases - an initial gradual but progressive forgetting phase followed by abrupt or catastrophic forgetting phase. Both gradual and catastrophic forgetting limit the usefulness of model editing methods at scale - the former making model editing less effective as multiple edits are made to the model while the latter caps the scalability of such model editing methods. Our analysis also highlights other key limitations of ROME and MEMIT at scale. With our work, we push for the development and evaluation of model editing methods keeping scalability in mind.
Model Editing at Scale leads to Gradual and Catastrophic Forgetting
Akshat Gupta, Anurag Rao, Gopala Anumanchipalli UC Berkeley akshat.gupta@berkeley.edu
1 Introduction
Editing knowledge in large language models (LLM) has recently emerged as a sought after capability for natural language processing (NLP) practitioners due to the huge retraining costs associated with LLMs. It is a known fact that LLMs memorize some part of the their training data (Radford et al., 2019; Carlini et al., 2022) in the form of stored knowledge. Model editing, sometimes also called knowledge editing [^1] In this paper, we will be using knowledge editing and model editing interchangeably, is the task of modifying existing knowledge or injecting new facts into the model Cohen et al. (2023), without updating the entire model. While knowledge in LLMs can be updated by continually pre-training models on new facts, this is not always viable due to the huge costs of training LLMs. Even then, LLMs are not perfect memorizers of training data Carlini et al. (2022) and will inevitably memorize facts incorrectly, thus requiring the capabilities of editing the incorrectly memorized facts. This is what makes model editing a very useful tool in the arsenal when working with LLMs.
Various methods have been proposed over the years to do so. Some of these methods (De Cao et al., 2021; Mitchell et al., 2021) require training a hypernetwork Chauhan et al. (2023) that generates the new weights for the model being edited. Other methods (Meng et al., 2022a, b) directly update specific parts of the model after identifying location of stored facts inside it (Geva et al., 2020). The success of knowledge editing is usually evaluated along three dimensions (Yao et al., 2023; De Cao et al., 2021; Meng et al., 2022a, b) - (i) reliability, which measures if the post-edit model accurately recalls the newly edited fact, (ii) generalization, which measures if the post-edit model accurately recalls the newly edited facts for different phrasings of the edited fact, and (iii) locality (also known as specificity (Meng et al., 2022a, b)), which measures if the edited fact changes unrelated or neighboring facts.
While many of these methods have shown promise Yao et al. (2023), recent work analyzing the after-effects of these editing methods have highlighted the shortcomings of these methods. Specifically, while some of these editing methods rank high on reliability, generalization and locality metrics (Yao et al., 2023; Mitchell et al., 2021, 2022; Meng et al., 2022a, b), the edited knowledge is not used consistently by the model. Cohen et al. (2023) propose a new evaluation system where the "ripple effects" or implications of an edited fact are evaluated. An example of such ripple effects would be - if an edited fact updates the president of a country to the new president, then prompting for the birthplace of the president should output the birthplace of the new president. Li et al. (2023) extend this by introduce the concept of "knowledge conflict" and additional edit types like reverse-edits and round-edits, thus evaluating the logical consistency of model editing in more complex scenarios.
Even though recent works are beginning to evaluate model editing more thoroughly, one very important aspect that is not given enough attention is making multiple edits on the same model. Most of prior works (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a; Cohen et al., 2023; Li et al., 2023) focus on evaluating model editing performance when only a single edit is made to the model and find that making multiple edits, either in a sequential or batched manner, leads to significant degradation of performance (Mitchell et al., 2021, 2022; Meng et al., 2022b). In an ideal realization of editing knowledge in LLMs, these methods will be used to updates thousands if not hundreds of thousands of facts. When multiple edits were made to the same model sequentially, Meng et al. (2022b) see a significant decrease in performance for all methods, starting as low as 10 edits. This brings to attention a very important question - do these methods scale?
With the popularity of in-context learning Dong et al. (2022) and retrieval augmented generation (RAG) methods Lewis et al. (2020), knowledge editing only starts to become a viable alternative when the number of memories being edited or injected compete with the context length of the model. This means that knowledge editing has almost no practical use when used to edit just one fact and only starts to become useful at scale. Which is why in this paper, we keep scalability at the center when evaluating model editing in LLMs. We make multiple edits to the same model sequentially and analyse the effects of these edits on the model as a function of number of edits. Specifically, we use two best performing knowledge editing methods - ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b) and make multiple edits on the same model. Along with the usual metrics for evaluating knowledge editing, we propose evaluating post-edit models on downstream tasks to make sure that model editing doesn’t hurt the regular functions of a model. We do that by evaluating the zero-shot performance of edited models on two downstream tasks from the GLUE benchmark Wang et al. (2018).
We find that knowledge edits made by ROME and MEMIT are not as localized as they are believed to be. New edits consistently bleed into other facts stored in the model, including the previously edited facts. We also find the model editing methods are prone to gradual and catastrophic forgetting. To the best of our knowledge, our work is the first to associate model editing methods with catastrophic forgetting. We define gradual forgetting as a progressive loss of ability of the model to perform its regular functions as the model is continuously modified through knowledge edits. This includes forgetting the ability to recall previously edited facts as well doing downstream tasks. Gradual forgetting is followed by sudden or catastrophic forgetting (Goodfellow et al., 2013; Kirkpatrick et al., 2017), where the model gets crippled due to a single update made to the model. We call such edits made to the model disabling edits. A disabling edit decapitates the model - new knowledge edits are no longer successful on the model, the model forgets all previously edited facts and is unable to perform any downstream tasks. We also analyze different properties of and reasons behind such disabling edits. With this paper we highlight some serious limitations of current model editing techniques, especially their lack of robustness when scaled and call for further research in developing scalable model editing methods.
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 1: This figure shows the editing proficiency of ROME on GPT-2XL using three random samples of the CounterFact dataset. The dotted line represents the metric averaged over a past window size of 5, whereas the solid lines represent the metric averaged over a past window size of 50.
Specifically, we make the following contributions in this work:
  • We extensively evaluate ROME and MEMIT at scale and highlight some of their key features and limitations.
  • We propose evaluating model editing methods on downstream task performance in addition to knowledge editing metrics. This becomes even more important when editing facts at scale.
  • We find that models edited using ROME and MEMIT continuously undergo forgetting of previously edited facts as well as the ability to do downstream tasks. This forgetting takes place in two phases - a gradual but continuous phase followed by catastrophic forgetting. These behaviours limit the scalability of model editing. To the best of our knowledge, our work is the first to associate model editing methods with catastrophic forgetting.
  • We discover disabling edits - these are single knowledge edits that cause the model to lose all function. We find that disabling edits are fundamental to ROME, while MEMIT is more robust to such single disabling edits.
2 Dataset and Methods
In this paper, we evaluate two most popular and best performing knowledge editing methods in literature - Rank-One Memory Editing (ROME) Meng et al. (2022a) and Mass-Editing Memory in a Transformer (MEMIT) Meng et al. (2022b). We evaluate these models on the CounterFact dataset Meng et al. (2022a), which contains 21,919 counterfactual statements. Each datapoint in the dataset is a triplet of the type (subject, relation, object). The dataset is created such that the target object is less likely than the original object in the relation triplet. This makes each datapoint in the dataset much harder to edit. In this preprint, we perform all experiments on GPT-2XL Radford et al. (2019) due to computational constraints. These experiments will be expanded to other models in future iterations of the paper.
3 Scaling ROME
In this section, we evaluate the performance of Rank-One Memory Editing (ROME) method Meng et al. (2022a) when multiple edits are made to the same model. These edits are made using the CounterFact dataset. To perform these experiments, we create three random subsets of 1000 examples from the CounterFact dataset and sequentially edit the model on the selected datapoints. We do so to find patterns that are independent of the effect of the order in which facts are edited. In this paper, we present three representative subsets that represent three different ways in which model editing dynamics proceed for the ROME method. There is no knowledge conflict Li et al. (2023) in any of these samples as the same subject is not edited twice in a subset.
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 2: This figure shows the percentage of facts forgotten as a function of number of edits made to GT-2XL using ROME on the same three samples as used in Figure 1. The percentage is calculated over the number of facts that were successfully edited up to given number of edits.
3.1 Editing Proficiency at Scale
[…]
Figure 1 shows the different metrics used for measuring editing proficiency of ROME on GPT-2XL as a function of the number of edits made to the model. The dotted lines represent average over a past window of size 5, which is the average of the given metric over 5 previous edits. The solid lines represents the average metric over a window size of 50. We find that the efficacy score, which measures the success of an edit, is almost 100% until a point where it begins to decline, sometimes very steeply. The three samples show that this point of steep decline can come as early as 100 edits, or as late as 800 edits made to the model  [^3] Experiments with other subsets have shown that this inflection point can occur at as early as 10 edits made to the model.. While ROME is successful at making edits to the model prior to this inflection point, its neighborhood accuracy still consistently declines as more edits are made to the model, indicating that the edits made start to bleed into other fact stored in the model, as more edits are made to the model. We will provide more evidence for this in later sections.
To summarize, we find that ROME is extremely proficient at successfully editing knowledge but only up to a certain point, after which, suddenly the method breaks down. Although even before that point is reached, the edited knowledge does begins to impact the unrelated knowledge stored in the model, as shown by the consistently declining neighborhood score.
3.2 Gradual and Catastrophic Forgetting
As new facts get added successfully to the model, is the model able to remember previously edited facts?
Figure 2 shows the number of previously correctly edited facts that get forgotten as a function of new edits made to the model. Firstly, we want to point the reader to clear spikes in all three graphs that occur at approximately the the same locations where the efficacy score in Figure 1 starts to decline. This indicates that the inflection point observed in Figure 1 has additional consequences other than just the loss of editability of the model..
Prior to this inflection point, Figure 2 shows an increasing relationship between the number of forgotten facts and the number of edits made to the model [^4] The graphs in figure 2 are evaluated after every 10 edits for computational reasons.. For example, for Sample 1, the model forgets approximately 30% of edited facts after 600 edits made to the model. Thus prior to the inflection point, we see a region where the model gradually forgets the previously edited facts. Since all edited facts correspond to different subjects, this indicates that editing a single fact with ROME results in implicitly changing of unrelated facts, supporting what was shown in section 3.1. After this region of gradual forgetting of facts, we reach an inflection point where we find that a catastrophically large number of facts are forgotten by the model. This is the same point where any further knowledge editing also suddenly starts to become unsuccessful (Figure 1).
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 3: This figure shows the downstream performance of editing GPT2-XL using ROME on SST2 and MRPC tasks.
This phenomenon of sudden forgetting of a huge number of facts is a realization of catastrophic forgetting in machine learning literature (Goodfellow et al., 2013; Kirkpatrick et al., 2017). Catastrophic forgetting is defined as the sudden loss of ability of a model to perform a prior task when it is further trained to perform a new task. The phenomenon observed in the above example is a perfect realization of how "catastrophic" or abrupt catastrophic forgetting can be, where it literally "forgets" an exploding number of facts with one gradient update. To the best of our knowledge, our work is the first to show that knowledge editing methods are also prone to catastrophic forgetting. But is catastrophic forgetting just limited to abruptly forgetting previously edited facts? In the next section, we show that it goes beyond just forgetting of previously edited facts.
3.3 Downstream Evaluation of Edited Models
One implicit feature expected out of all model editing methods is that as a fact is edited or inserted into the model memory, it should not effect the model’s ability to perform its regular functions. This means that knowledge editing should not effect the model’s ability to perform common NLP tasks which the model is used for. We call this an implicit assumption because to the best of our knowledge, none of the previous works try to directly measure the effect of model editing on downstream tasks. Meng et al. (2022a, b) try to capture this effect in some form by calculating the generation entropy, which measures weighted average of the bi-gram and tri-gram entropy of text generated by the post-edit model. The idea is to measure model degradation when generating text using a post-edit model. Ideally, we want the generation entropy of cases with lots of repetitions (a common failure case of model editing (Meng et al., 2022a, b)) to be lower than coherently generated text. We find that this metric is unable to directly quantify the degradation of model ability as multiple edits are made to the model. We explain this in more detail and give examples of failures of generation fluency metric in appendix A.1.
We take a more direct approach to quantify model degradation by measuring the performance of the post-edit model on common downstream NLP tasks. We chose two tasks from the popular GLUE benchmark Wang et al. (2018) - sentiment analysis (SST2) Socher et al. (2013) and paraphrase detection (MRPC) Dolan and Brockett (2005) for this. Both are binary classification tasks and we use a balanced subset of 200 test examples and evaluate the models using the accuracy metric. We evaluate the downstream model performance on these tasks as a function of number of edits made to the model in a few-shot setting. The in-context examples and prompt template used for evaluation are presented in Tables 2 and 3.
Figure 3 depicts the effect of model editing on downstream tasks as a function of the number of edits made to the model. We again see two regions of model degradation. Initially, there is a gradual decline in downstream performance of the model for both SST-2 and MRPC tasks with an increasing number of edits made to the model. We then see an inflection point with an abrupt loss of ability of the model to perform either tasks, which coincides with the point of sudden decrease in editability of the model (section 3.1) and a catastrophic increase in forgetting previously edited facts (section 3.2). Additionally, we want to point the user to Sample 1 in Figure 3 which contains the maximum number of edits made to the model before the inflection point. We see that the model is no longer able to perform paraphrase detection long before the inflection point, showing that as the model is edited multiple times, it continuously loses its ability to perform downstream functions and may be unusable for certain downstream tasks even before the inflection point.
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 4: This figure shows the normalized L2 distance between the edited layer weights and its original weights as a function of the number of edits made to the model. The editing algorithm is ROME and the model being edited is GPT2-XL.
While the inflection point is a big concern for model editing methods, there is also a gradual decrease of general ability of the model even prior to that point, which is only visible if the model is evaluated on a downstream task. This shows the usefulness of evaluation of model editing methods on downstream tasks, which we urge the research community to adopt along with other knowledge editing metrics. We define the first region where the model progressively loses its ability to do prior tasks (like recalling previously edited facts or performing downstream tasks) as gradual forgetting, juxtaposing it with catastrophic forgetting. Note that forgetting here does not just refer to forgetting previously edited facts but a general loss of ability to perform a certain function. We find that sequential editing of a model always leads to these two phases of forgetting - gradual forgetting and catastrophic forgetting.
3.4 The Source of Forgetting
So far, we’ve seen that sequentially editing of multiple facts in LLMs leads to the model gradually forgetting previously edited facts and losing the ability to be useful for downstream tasks. This is followed by an abrupt inflection point, which not only leads to a forgetting almost all previously edited facts, but also a complete loss of model ability to perform regular NLP tasks, thus rendering the model useless. This inflection point is a fundamental feature of ROME which can be seen across all samples of multiple sequential edits, and is a realization of extreme catastrophic forgetting. But is this point an outcome of continuous editing of the model or the result of a single edit to the model? What is the reason behind these two phases of forgetting? In this section, we answer these question in more detail.
Model editing methods are designed with the objective of editing or inserting specific facts stored inside the model without changing all the weights of the model. A precursor to be able to do this is localizing a fact down to specific neurons inside a model and then only changing the weights of the identified neurons. The ROME method is built on the assumption that a fact can be changed by changing the weights of any one out of a set of knowledge-storing layers of a model while keeping the rest of the model the same, which is showed to work empirically and backed by causal tracing experiments Meng et al. (2022a). Each time we make such edits, the edited layer becomes slightly different from its original version. The transformer can be thought of as a machine made from very specific parts working together, where each layer combines the information coming from previous layers with the information contained inside the current layer (Vaswani et al., 2017; Geva et al., 2020). To be able to do this, each layer must be able to understand the signal coming from the previous layers. In simpler words, there is a notion of compatibility between the different layers of the transformer when they are trained together. As we edit one specific layer of the model continuously while keeping the rest of the model constant, we are constantly changing one part of the model while keeping the remaining part the same. Such a procedure is bound to reach a point where the layer that is changed becomes so different from its original version that this compatibility is destroyed and the other parts of the transformer are unable to makes sense of the incoming signal from the layer being edited.
(a) Editing Proficiency
(b) Forgetting
(c) Downstream Performance
Figure 5: This figure shows the editing proficiency of MEMIT on GPT-2XL for Sample 1 over 5000 sequential edits made to the model.
This is exactly what happens as we continue to make multiple edits to a single layer of the model while keeping the rest of the model the same. This can be seen in Figure 4. Figure 4 shows the normalized [^5] We first take the L2 norm between original and post-edit weights of the layer being edited, and then normalize it by number of neurons in the layer. L2 distance between the weights of the edited layer and the original weights of the layer as a function the number of edits made to the model. We find that as more edits are made to the model, the distance between the original and edited layer continuously increases until it suddenly explodes. This is the point where the edited layer becomes incompatible with the rest of the model. At this point, the model breaks down and catastrophically forgets previously learnt facts, loses its ability to do downstream tasks and its ability to be correct by model editing methods. The gradual increase in distance between the original weight and new weights leads to the gradual region of forgetting, whereas the spike in the distance with a single updates leads to catastrophic forgetting.
We can clearly see that the spikes in Figure 4 exactly coincide with the spikes in Figures 2 and Figure 3. This also explains the multiple spikes in Figure 2. As the model reaches the inflection point, the model editability also goes down but ROME is still able to edit some facts with limited accuracy. At this point, the model is over-fitting on the edited facts as it is only able to retrieve the recently edited facts and has completely forgotten previously edited knowledge and the ability to do downstream tasks. But as the model is edited further, we again see a spike in distance from the weights of the model at the previous spike, and the model again goes through catastrophic forgetting of the recently learnt facts.
Finally, we take a deeper look at the specific edits that cause the inflection point. We call these edits disabling edits, as they disable the model and make it unusable for downstream tasks. Are these disabling edits a result of continuous sequential editing that accumulates over time or of one specific fact that is especially hard for a model to learn? We find that when we edit the facts corresponding to the disabling edits as the first edit made to the model, the model is still left completely disabled, that is, it is still unable to perform any downstream tasks, and the normalized distance of the layer weights from the original weights is comparable to the distances seen around the spikes. This shows that the disabling edits in ROME are not a result of continuous sequential editing of the model, but are a fundamental limitation of ROME. We can describe disabling edits as facts that ROME is unable to successfully edit without crippling the model. This also disproves the fundamental hypothesis behind ROME that knowledge in LLMs can be edited by changing the weights of a single layer inside the model.
We also want to point out that while existence of such disabling edits are a fundamental limitation of ROME, it can also be used as a feature. If we can test the editability of a fact prior to editing by checking if it is a disabling fact, we can possibly prevent catastrophic forgetting when using ROME. Such explorations are beyond the scope of this paper and should be part of future research.
4 Scaling MEMIT
In this section, we evaluate the performance of the Mass-Editing Memory in a Transformer (MEMIT) method Meng et al. (2022b) when multiple sequential edits are made to the same model. We will follow the same procedure as followed in section 3, first evaluating the editing proficiency, gradual and catastrophic forgetting and loss of performance on downstream tasks. We perform sequential editing on GPT2-XL using a random subset of 5000 examples from the CounterFact dataset when using MEMIT. This subset is a continuation of sample 1 in section 3. The behavior of MEMIT when sequentially editing GPT2-XL is shown in Figure 5.
Figure 4(a) shows the editing proficiency of MEMIT as a function of the number of edits made to the model. Note that here we edit one fact at a time for MEMIT. While MEMIT is able to make batched edits, we leave that analysis for future work. The dotted lines show a window size of 5 previous edits, whereas solid lines show a window size of 50 previous edits, same as in Figure 1. We see that the efficacy score for MEMIT is not as high as ROME. This means that knowledge edits made via MEMIT are not always as successful as in ROME. We also see a continuous decline of neighborhood score for MEMIT as seen for ROME, showing that editing facts possibly start affecting other facts stored in the model.
Figure 4(b) shows the percentage of successfully edited facts that get forgotten as new facts are edited using MEMIT. We again begin to see two phases of forgetting - gradual and catastrophic. Forgetting is evaluated in Figure 4(b) after every 500 edits beyond 2000 edits for computational reasons, which is why the plot looks spiky afterwards. For this reason, we want to clearly point out the region of gradual forgetting to the reader, which goes on until approximately 4000 edits, which is followed by the region of catastrophic forgetting after which the model forgets all previously learnt facts. When compared with the gradual forgetting region of Figure 2, we find that MEMIT forgets fewer previously edited facts when compared to ROME. This is seen across multiple samples of edits made using MEMIT, although we show only one due to space constraints. As a comparison, model edited with ROME forgets approximately 30% facts after 600 edits in sample 1 (Figure 1(a)), whereas when editing the same sample with examples in the exact same order, MEMIT forgets approximately 10% previously edited facts. Additionally, we don’t encounter an inflection point for a larger number of edits when compared to ROME. This is true for multiple samples edited using MEMIT. This shows that although MEMIT is prone to gradual forgetting, it has a longer gradual forgetting phase, thus making it more stable when making a larger number of sequential edits to the model.
Figure 6: This figure shows the normalized L2 distance between the edited layer weights and its original weights as a function of the number of edits made to the model. The editing algorithm is MEMIT and the model being edited is GPT2-XL.
Finally, the effect of the number of edits on the downstream performance of the model can be seen in Figure 4(c). We see that the model loses its ability to do tasks gradually as more number of edits are made to the model. We again see two regions of forgetting - a gradual loss in ability and then a sudden drop in the model’s ability to do downstream tasks. The model loses the ability to do paraphrase detection long before the point of catastrophic forgetting, thus showcasing that the model can lose its ability of performing certain downstream tasks even before a disabling edit cripples the model.
Figure 6 shows the distance of the edited layers from the respective original layers. We find that the distance from respective original layer increases gradually until 4000 edits, which is also the region of gradual forgetting of previously edited facts and downstream ability. After this, we find a sudden spike in the distance between the edited layers and the original layers, which coincides with the points of catastrophic forgetting as seen in previous plots. We also evaluate if disabling edits are fundamental to MEMIT, and find that if the fact that disables the model after a sequence of edit is edited first in the model, it does not lead to catastrophic forgetting. This means that in MEMIT, the effect of disabling edits is normalized. This supports the hypothesis behind MEMIT that facts are possibly stored across multiple layers and to be able to successfully edit facts, we need to edit the weights across each of these fact storing layers.
5 Limitations of ROME and MEMIT at Scale
We find that ROME and MEMIT have significant functional differences. ROME has almost 100% efficacy when editing facts although the model ends up forgetting a larger number of previously edited facts as multiple edits are made sequentially. MEMIT on the other hand is unable to successfully edit all facts, but fewer facts get forgotten as multiple edits are made to the model. Both ROME and MEMIT have low specificity and edited facts consistently bleed into other knowledge stored inside the model, especially with larger number of edits made to the model. MEMIT is also more robust to a single de-stabilizing edit when compared to ROME. We conjecture this is because a fact is stored within multiple layers of a model, and editing the weights of a single layer to edit facts can lead to instability in the model. Additionally, we see that both ROME and MEMIT are prone to forgetting in a machine learning sense, losing the ability to recall previously edited facts, to perform downstream function and the ability to be successfuly edited by these algorithms. This makes these algorithms hard to scale and not usable for practical model editing purposes.
6 Conclusion
In this paper we analyze two popular model editing techniques - ROME and MEMIT, at scale. We use these methods to make multiple sequential edits to the same model and find that they fail in multiple ways. We discover that the model being edited with these techniques undergoes forgetting of previous knowledge and skills in two phases. Initially, it gradually forgets previously edited facts and loses the ability to do downstream tasks, a phase which we call gradual forgetting. After that, the model abruptly loses all coherence and function including the ability to recall previously edited facts and perform downstream tasks, which is a realization of catastrophic forgetting. We also find that the source of these two phases of forgetting is that the layers being edited with these methods slowly drift away from their original weight values, thus becoming incompatible with the rest of the model.
Practical use of model editing requires us to be able to make multiple sequential edits to a model. The above findings indicate serious limitations of current state-of-the-art model editing methods at scale. We find that these model editing methods, like other fine tuning techniques, are prone to catastrophic forgetting. To be able to scale such methods, we not only need to have high efficacy, specificity and generalization, but we also need these methods to preserve the model’s existing abilities. Thus, it becomes essential to evaluate post-edit models on downstream tasks. Also, as multiple facts are edited inside a model, the model should remember previously edited facts. With this paper, we call for an improved evaluation of model editing techniques, including evaluating model performance on downstream tasks and ability to recall previously edited facts.
Finally, we want to stress upon the implications of the two phases of forgetting discovered in this paper. Gradual forgetting makes model editing techniques become increasingly less effective as we sequentially edit facts, and hence limits their usefulness at scale. While catastrophic forgetting, which renders the model practically useless, caps the extent to which we can scale these methods. Thus we need to create model editing techniques that can counteract both gradual forgetting and catastrophic forgetting when scaled.
Appendix A Appendix
A.1 Generation Fluency Examples
We present the limitations of the generation entropy metric as proposed by (Meng et al., 2022a, b) in Table 1. The metric of generation entropy measures the weighted average of the bi-gram and tri-gram entropy of generated text distribution post editing. The aim of the metric is to be able to quantify model degradation while making multiple knowledge edits to the model. The specifically aims to quantify a common failure mode in knowledge editing when the post-edit model generates a lot of repeated words. In this metric, the entropy of the generated text distribution should be lower if there is a lot of repetition in generated text, and higher if there is more variety, thus higher is better for this metric.
We present examples of generated text when an edit is made successfully and unsuccessfully in Table 1. The first row in Table 1 represents a successful knowledge edit, where a model is able to recall the edited fact post-editing successfully. The second row represents an unsuccessful edit, where the model is not able to recall the edited fact post-editing. Also, the generated text contains a lot of repetitions and spurious symbols and the model has degenerated syntactically and semantically. Yet, the generation entropy for the second row is higher than the first row. This example shows the first flaw in the generation entropy metric, where it is not able to differentiate between text with lots of repetitions and incorrect grammar. The third row in Table 1 shows another example of successful model edit, where the generated text contains semantically sound English. The generation entropy score for the third row is identical to the second row, which clearly shows that the metric is also unable to quantify and differentiate between grammatically sound English and gibberish, as it gives both row two and three the exact same generation entropy value. Because of this, we conclude that the generation entropy metric is not an informative metric to measure model degradation.
Edit Success
Generated Text
Generation Entropy
True
One can get to Plana Peak by navigating to the southern side of the mountain. To the left of the peak is the path to the north of the mountain, which leads to the entrance of the mountain’s south side. The path to the north is blocked by a rock. To the right of the rock is the path to the north. The mountain is surrounded by a large cliff. To the left is the entrance of the mountain’s northern side. The path to the left is blocked by a
5.96
False
Toronto’s twin city has famous tourist attractions including the is, and Syrian aid,,, and aid. And and, and and, and, which, is, [ily, and is and,, and and and [ily was, no, and, and,,, and. no aidily, and Syriaily, do, no doubt de [ and which no Syrian is do a, no., and the, is and is, and, and and,, [ aid, and
6.08
True
Brut y Tywysogion is full of people who speak a language other than their own. The French, for example, speak French, the Germans speak German, the Italians speak Italian, and so on. The French and Germans speak French, the Italians speak Italian, the Americans speak English, the Japanese speak Japanese, and so on. The English speak English, the Americans speak English, the Japanese speak Japanese, and so on and so forth. The Japanese
6.08
Table 1: Text generated by post-edit model for ROME and corresponding generation entropy scores.
Task
Few-Shot Prompt
SST-2
Review : excruciatingly unfunny and pitifully unromantic
Sentiment : negative
Review : rich veins of funny stuff in this movie
Sentiment : positive
Review : by far the worst movie of the year
Sentiment : negative
Review : fashioning an engrossing entertainment out
Sentiment : positive
Review : INPUT SENTENCE
Sentiment :
Table 2: Few-shot template used to measure downstream model performance for SST-2 task.
Task
Few-Shot Prompt
MRPC
Are the sentences paraphrases of each other.
Sentence 1: Mr McDevitt has been granted control of three crucial aspects of policing in the Solomons.
Sentence 2: Mr McDevitt has been granted control of three aspects of policing by Commissioner William Morrell.
Answer: No
Are the sentences paraphrases of each other.
Sentence 1: The notification was first reported Friday by MSNBC.
Sentence 2: MSNBC.com first reported the CIA request on Friday.
Answer: Yes
Are the sentences paraphrases of each other.
Sentence 1: In 2002, Linksys overtook Cisco Systems as the leading wireless equipment vendor, accounting for 14.1 percent of revenue.
Sentence 2: Rolfe said Linksys overtook Cisco Systems last year as the leading supplier of WLAN equipment.
Answer: No
Are the sentences paraphrases of each other.
Sentence 1: "The anticipated global sales improvement in the month of June did not materialize", said Chief Financial Officer Robert Rivet.
Sentence 2: "The anticipated global sales improvement in the month of June did not materialize as we had anticipated", the company said.
Answer: Yes
Are the sentences paraphrases of each other.
Sentence 1: That compared with \$ 35.18 million, or 24 cents per share, in the year-ago period.
Sentence 2: Earnings were affected by a non-recurring \$8 million tax benefit in the year-ago period.
Answer: No
Are the sentences paraphrases of each other.
Sentence 1: They had published an advertisement on the Internet on June 10, offering the cargo for sale, he added.
Sentence 2: On June 10, the ship’s owners had published an advertisement on the Internet, offering the explosives for sale.
Answer: Yes
Table 3: Few-shot template used to measure downstream model performance for MRPC task.